If you have 7.0 moles O2, how many moles of H2 are needed for the reaction?
1
Expert's answer
2020-10-05T14:27:32-0400

The reaction between H2 and O2 is as follows

H2 + (1/2)O2 = H2O

So here according to the reaction 0.5 mol O2 require 1 mol H2.

So for 7 mol O2 , H2 require

=( 7/0.5) mol

= 7×2 = 14 mol (ans)